If the random variable x follows a normal distribution with a mean of 18 and a standard deviation of 5, find the value of k such that

1-p(X>K)=0.2578

2-p(X<K)=0.2578


1
Expert's answer
2021-12-09T03:02:44-0500

"\mu=18 \\

\sigma = 5"

1.

"P(X>K) = 0.2578 \\

1 -P(X<K) =0.2578 \\

P(X<K) = 0.7422 \\

P(Z< \frac{K-18}{5}) = 0.7422 \\

\frac{K-18}{5} = 0.6505 \\

K -18 = 5 \times 0.6505 \\

K = 18 +3.2525 = 21.2525"

2.

"P(X<K)=0.2578 \\

P(Z< \frac{K-18}{5}) = 0.2578 \\

\frac{K-18}{5} = -0.65 \\

K-18 = 5 \times (-0.65) \\

K = 18 -3.25 = 14.75"